40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 34     Entry time: Wed Oct 31 08:33:54 2007
Author: rana 
Type: Problem Fixed 
Category: SUS 
Subject: Vent measurements 
There was a power outage during the day yesterday; whoever was around should post something here about the
exact times. Andrey and David and Tobin got the computers back up - there were some hiccups which you can
read about in David's forthcoming elog entry.

We restarted a few of the locking scripts on op340m: FSSSlowServo, MCautolocker. Along with the updates
to the cold restart procedures we have to put an entry in there for op340m and a list of what scripts
to restart.

David tuned up the FSS Slow PID parameters a little; he and Andrey will log some entry about the proper
PID recipe very soon. We tested the new settings and the step response looks good.

We got the MC locking with no fuss. The 5.6 EQ in San Francisco tripped all of the watchdogs and I upped
the trip levels to keep them OK. We should hound Rob relentlessly to put the watchdog rampdown.pl into
the crontab for op340m.
ELOG V3.1.3-